A population is estimated to have a standard deviation of 12. We want to estimate the population mean within 2, with a 95% level of confidence.

How large a sample is required? (Round up your answer to the next whole number.)
Answer #1

Given, x=0.0s, 6=12, 4=2 A = 20.025 * 5 ie n= 20.028 * ² = (1-959964)** 123 = 138.2925197 In = 139] .. Required sample size =